[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Feb 02, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Feb 2 19:48:17 UTC 2017


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
Late morning rounds.

Sheephead workstation is operational. NCO replaced the video driver to fix  
the problem encountered on mid shift.

NCF continues investigate the problems sending text products in GFE. GFE is  
still sluggish.
HSU:
Tuna is still down after the red hat upgrade.
Other items:
We continue to clean up the problems with the storm surge NHA domain after  
the upgrade.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170202/639b9935/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list